The center of the circle (x+1)^2 + y^2 = 4
Please put the solution.. thank you for being so nice and helpful tutors..^^

Answer by mgmoeab(37)

The center of the circle (x+1)^2 + y^2 = 4
Remember the equation of a circle with center (h, k) is given by:
(x-x1)^2 + (y - y1)^2 = r^2 were the center is (x1, y1)
***
in this case:
(x+1)^2 + y^2 = 4
***
which can be written as:
(x-(-1))^2 + (y - 0)^2 = 2^2 were the center would be (-1, 0) and a radius of 2.
hope it helps!